OpenStudy (anonymous):

Find the distance between the points (2, -3) and (5, -4)

OpenStudy (rane):

use this formula

OpenStudy (anonymous):

$$\sqrt{58}$$

OpenStudy (rane):

\[ y = \sqrt{(y2- y1)^2 + (x1 - x2)^2}\]

OpenStudy (rane):

where (2.-3) 2 is x1 nd -3 is y1 (5,-4), 5 is x2 nd -4 is y2

OpenStudy (anonymous):

Basically your forming a right triangle with legs equal to the difference of each corresponding coordinate in the two tuple, and using the Pythagorean theorem to solve for the distance between them.

OpenStudy (anonymous):

sqrt58 is correct answer?

OpenStudy (rane):

@Jack174 there is a Pacific formula to find the distance between 2 points. pls dont confuse thhat person with using the pythagorus theorem

OpenStudy (anonymous):

i don't know

OpenStudy (rane):

use the formula that i have givn u

OpenStudy (rane):

do u want me to do it f u didnt understand it

OpenStudy (anonymous):

yes please

OpenStudy (rane):

ok

OpenStudy (dumbcow):

the formula comes right from the pythagorean thm, so to understand what you are doing conceptually the pythagorean thm approach is better than memorizing a formula

OpenStudy (anonymous):

?

OpenStudy (anonymous):

is the distance 10

OpenStudy (anonymous):

do i need to put sqrt for answer or just number?

OpenStudy (anonymous):

(5-2) , (-4--3) = (3,-1)

OpenStudy (rane):